[Hallc_running] [New Logentry] RC Daily Update

mpaolone at jlab.org mpaolone at jlab.org
Sat Mar 2 14:00:08 EST 2019

Logentry Text:
Hi Everyone,

No beam since last night (~22:00).  While MCC was tuning 1L25, a vacuum valve would not open.  After inspection by the vacuum group, a damaged O-ring was noticed, and valve open/close mechanism had issues. (see elog for some more details:  https://logbooks.jlab.org/entry/3662073).   As of this post (~14:00), the valve seems to be open, so tuning of 1L25 can continue.  After it is tuned (or bypassed), MCC will check if beam transports beyond 1L25.  If so, we are back in business. Latest estimate is: we will know by 8pm tonight.

There are no major issues concerning the hall, and the run plan is to stay at the same kinematic setting (kin-4) through the end of the experiment.  

I see no need for a 4:30 meeting today, but I will update the log book when I get a new estimate of when we'll get beam back.  If you have questions, you can call me on the RC cell.



This is a plain text email for clients that cannot display HTML.  The full logentry can be found online at https://logbooks.jlab.org/entry/3662082
-------------- next part --------------
An HTML attachment was scrubbed...
URL: <https://mailman.jlab.org/pipermail/hallc_running/attachments/20190302/6c7ed0f0/attachment.html>

More information about the Hallc_running mailing list